# Impact of in-situ nuclear networks and atomic opacities on neutron star merger ejecta dynamics, nucleosynthesis, and kilonovae

### v1

@@ 3.3 Effects of thermalization treatment
The thermalization prescription directly enters the nuclear term in Eq. (1). However, nuclear heating remains a subdominant contribution to the dynamics for most of the ejecta evolution. We show this behavior in Fig 8, which displays the instantaneous average contributions to the internal energy of the fluid for the BLh profile under different thermalization and opacity prescriptions. On the timescales relevant for the nucleosynthesis, the effects of the different thermalization schemes on the temperature and density evolution are almost negligible. As a result, for the BLh profile, the final nuclear yields agree within roughly $\sim 10\%$. The isotopic evolution is similar for both thermalization schemes. Neutron freeze-out occurs slightly earlier (by $\sim 200$ ms) in the simple thermalization case, with no qualitative change in the rest of the isotopic evolution.
@@
At early times, the thermalization efficiency is close to unity for all decay products. Any constant, fractional thermalization factor therefore underestimates the nuclear contribution to the internal energy of the ejecta. The resulting lower temperature of the photosphere explains the dimmer and redder light curves produced by the simple thermalization scheme at $t\!\sim$ hours. The two models cross between roughly $\sim\!6$ hours and $\sim\!1$ day, when the effective averaged thermalization factor from the online calculations drops to $\langle f_{\text{th}}\rangle\sim 0.5$.
@@ 3.4 Effects of opacity models
The (optical) opacity enters the luminosity term in Eq. (1), and can influence the ejecta dynamics. Consequently, a different opacity model could also indirectly affect the composition evolution. However, since radiation contributes only a subdominant fraction of the energy balance (see Fig. 8), a detailed multi-group treatment of the opacity is not necessary to accurately capture the hydrodynamics. The radiation term becomes relevant only at $t\gtrsim 1$ days, when the expansion is already fully homologous and changes in the composition of the outflow are driven only by nuclear decays. An order-of-magnitude estimate of the gray opacity is therefore sufficient to correctly reproduce the ejecta dynamics and perform reliable nucleosynthesis calculations. The relative difference between the abundances predicted by different opacity models always remain below a few $0.1\%$.
@@
Figure 11: Evolution of the gray opacity $\kappa_{\textrm{gray}}$ for a sampling of fluid elements across all angles and depths, from simulations using different opacity models. Top row: Runs using the analytical expression from Wu et al. (2022) (left) or the prescription from Just et al. (2022) (right). Central row: Simulations with Rosseland (left) or Planck (right) gray opacities computed from the LANL data for neodymium as representative of the lanthanides. Bottom row: Rosseland mean opacity from the multi-group models using only neodymium to represent lanthanides (left) or also uranium as representative of actinides (right). The color bar is cut at $\kappa_{\textrm{gray}}=10^{-2}$ cm2g-1 for easier visualization. For each simulation, the top and bottom sub-panels show the trajectories of the fluid elements in the density-temperature and density-cumulative lanthanides mass fraction ($X_{\text{lan}}$) planes, respectively. Markers indicate different orders of magnitudes of $X_{\text{lan}}$, with circles for $X_{\text{lan}}<5\times 10^{-5}$, downwards triangles for $10^{-5}<X_{\text{lan}}<10^{-3}$, upwards triangles for $10^{-3}<X_{\text{lan}}<5\times 10^{-2}$, and stars for $X_{\text{lan}}>5\times 10^{-2}$. For the simulations using only neodymium data, the background shows the single-element mean opacity $\kappa_{\textrm{el}}$ calculated from the LANL tables (Rosseland for GK and Nd300, Planck for GKPl). Similarly, the uranium Rosseland mean opacity is shown in the NdU300 panel.
The same reasoning applies when comparing the multi-group calculations with the J[22] model, employed for the JK run. Although this prescription adapts better to the multi-group results based on neodymium (and uranium) data (simulations Nd300 and NdU300), it still overestimates the opacity in the (equatorial) lanthanide-active ($X_{\text{lan}}\gtrsim 10^{-5}$) regions at temperatures around $T\sim 10^{3}$ K.
@@
The non-analytical models exhibit non-monotonic opacity evolution in the lanthanide-rich elements. This behavior arises either from the opacity peak around $T\sim 5\times 10^{3}$ K in the LANL-R and LANL cases, or from the activation of the atomic data at $T\simeq 6\times 10^{4}$ K in the LANL-P model (see Fig. 11). As a result, the photosphere is observed to bounce towards the outer regions of the ejecta at $t\sim 1$ hour in the GKPl simulation, and later at $t\sim 5$ hours in the GK, Nd300, and NdU300 runs. The photosphere location predicted by the multi-group calculations always lies between those obtained with LANL-R and LANL-P, which are known to respectively under- and overestimate the effective gray opacity.
@@
In Fig. 13, we examine the impact of the opacity model on the kilonova light curves. The colder photosphere predicted by W[22] with respect to the LANL model results in a suppressed emission at $t\lesssim 1$ hour, with a stronger suppression at higher frequencies that makes the early-time signal appear redder. Nevertheless, the photospheric luminosity and the optical-band magnitudes remain broadly comparable across models, as the larger photospheric radius partially compensates for the lower emissivity of the colder photosphere. At later times, the artificially high constant gray opacity causes the ejecta to become completely optically thin only at comparatively late stages. Therefore, the photospheric contribution to the luminosity persists for longer, thus delaying the decay of the infrared emission at $t\gtrsim 1$ day.
@@
Figure 13: Predicted AB apparent magnitudes in the Gemini $u$, $g$, $i$ and $K_{s}$ bands for an observer at a polar angle of $30$ degrees and a distance of 40 Mpc. Comparison between a series of opacity models applied to the BLh profile. Left: Frequency- and composition-dependent opacities based on atomic calculations for neodymium and uranium (solid lines), versus the Just et al. (2022) prescription (dashed) and the simple Wu et al. (2022) opacity (dotted). Center: Opacities based on the neodymium atomic calculations, multi-group (solid lines) versus gray models (Rosseland or Planck, dashed and dotted). Right: Frequency-dependent LANL opacities including uranium (solid lines) or only neodymium (dashed) compared against the analytical Just et al. (2022) model (dotted).
@@
As discussed above, J[22] performs better than W[22], but still overestimates the opacity of the lanthanide-rich ejecta. As a result, the light curves from J[22] more closely resemble those produced with the LANL models than those from W[22], but they still exhibit a delayed shut-off of the photospheric emission and thus a more prolonged signal.
@@
The central panel of Fig. 13 compares simulations using gray (GK and GKPl) and multi-group (Nd300) opacities based on neodymium atomic calculations. As expected, the Rosseland mean opacity performs better in the light-curve calculations at early times, when the ejecta is still optically thick. At later times, the Planck opacity provides a better approximation for the kilonova emission, especially in the higher-frequency bands. The two gray approximations bracket the actual effective opacity, and therefore generally bracket the multi-group light curves across all filters over most of the evolution. In particular, the Planck mean misses the early blue/UV peak at $t\sim 0.1$ days, whereas the Rosseland opacity predicts too bright and delayed magnitude peaks, especially in the higher-frequency filters. At late times ($t\gtrsim 10$ days), the Planck opacity performs worse again because of the delayed recession of the photosphere in a more opaque material (see Fig. 12). This effect dominates over the direct suppression of the emergent flux (e.g. Eqs. (3) and (19)), leading to an inversion of the hierarchy of the models in the red filter at $t\gtrsim 5$ days.
@@
The right panel of Fig. 13 shows that J[22] reproduces multi-group results based on neodymium data more accurately than our gray, atomic-physics-based models. This result highlights the necessity of employing frequency-dependent transport when using detailed atomic opacities, as compressing the information from line-resolved calculations into effective gray averages can be misleading to the point of being outperformed by analytical fitting models.
@@
The original work of Kasen et al. (2017) employed atomic calculations only up to the lanthanides, which at least partially explains the disagreement with our most sophisticated opacity model that also includes uranium (NdU300). Fontes et al. (2020) demonstrated differences between semi-relativistic and fully-relativistic neodymium opacities. Additional improvements in opacity modeling may be achieved by replacing theoretically calculated energy levels with NIST-calibrated values. Incorporating these calibrated energies may affects the corresponding light curves and spectra at late times.
@@
When only neodymium is incorporated, the effective opacities predicted by the frequency-dependent simulations lie between the Rosseland and Planck mean values, but are qualitatively closer to the latter (see Fig. 11). Correspondingly, the J[22] approximation produces light curves more similar to LANL-P, particularly at late times, when the ejecta becomes optically thin. However, the analytical prescription fails to capture the drop in effective opacity for the $r$-process fluid elements at $T\sim 2-3\times 10^{4}$ K. As a result, it misses the luminosity peak at $t\sim 2-3$ hours, caused by regions of the outflow that are hotter than their surroundings due to early ($t\sim 1$ s) decays of freshly synthesized $r$-process elements (Bernuzzi et al., 2025). This missing peak, already visible in the original publication from Just et al. (2022), is more relevant for trajectories with higher lanthanide fractions.
@@ 3.5 Impact of a polar jet
As already commented in Magistrelli et al. (2024), the impact of the jet on the globally averaged nucleosynthesis is negligible. Figure 14 shows the nuclear yields from the polar regions ($\theta\lesssim 15$ degrees) of the BLh ejecta in the structured-jet and no-jet configurations. Because this angular cut excludes the neutron-rich equatorial outflow, the resulting nucleosynthesis is dominated by the production of elements up to the first peak. The nuclear yields are essentially unchanged by the introduction of the jet, with typical discrepancies of the order of few tens of percent or less.
@@ 3.6 Best model comparison
In this section, we summarize the combined effects of all the improvements that we implemented on top of the code from Wu et al. (2022). We compare our most advanced setup (NdU300), featuring the in-situ NN, detailed thermalization, multi-group radiative transfer, and atomic-physics based opacities, against the results obtained using the 2D ray-by-ray extension of original version of the code (Apr2_BLh). For the latter, we also implement the generalized fits for the nuclear power discussed in Sec. 2.3.
As already mentioned in the previous sections, the improved thermalization and opacity prescriptions do not significantly affect the composition evolution of the ejecta. Slightly larger discrepancies appear in the nucleosynthesis yields and abundances evolution with respect to those described in Sec. 3.1, but the physical considerations remain the same. The introduction of the in-situ NN is the dominant source of variation in the nucleosynthesis results, and a detailed description of the hydrodynamic evolution of the ejecta is essential for reliable nucleosynthesis predictions. In particular, assuming homologous expansion is not safe on $r$-process timescales, leading to deviations greater than unity for most mass numbers and a shifted third $r$-process peaks (see the green and red curves in the bottom panels of Fig. 3). On the other hand, calculating the nuclear power on the fly and including its back-reaction on the hydrodynamics has a comparatively minor impact, typically at the $\lesssim 20\%$ level (see the orange lines in the same panels). For the purpose of studying element formation, the hydrodynamics of the ejecta can therefore be reproduced well enough by performing simulations with accurately-calibrated nuclear power fits. A post-processing NN can then recover the results of the online calculations within a few $10\%$, provided that the set of tracers is sufficiently large to properly resolve both the dynamics and the initial composition distribution of the ejecta.
@@
Figure 16: Late-time ($t\simeq 3$ days) ratios between the mass-weighted histograms of the density, temperature and instantaneous thermalized heating rate obtained from the BLh runs using the in-situ NN and our most advance thermalization and opacity prescriptions (NdU300), or from the 2D ray-by-ray extension of the configuration of Wu et al. (2022) (Apr2_BLh). The left [right] column shows the ratios as functions of the polar angle [initial electron fraction]. Bins containing less than $10^{-5}M_{\odot}$ in both simulations are plotted with increasing transparency for decreasing mass content (considering the run with most mass in that bin). Thus, the color encodes the value of the ratio in $\Delta M$, while the saturation reflects the highest between the two original values $\Delta M$ (cf. Fig. 5)
@@
The differences in temperature evolution and thermalized heating rates have an immediate impact on the predicted light curves. Figure 17 displays the kilonovae associated with the Apr2_BLh and NdU300 simulations, both showing the characteristic blue/UV peak at $t\sim 1$ hour and the later red emission. At early times ($t\lesssim 0.1$ day), the average thermalization efficiency is comparable between the two simulations, and higher in the DT08 scheme. At the same time, the multi-group opacities drive a faster recession of the photosphere. Both effects contribute to a hotter photosphere than in the simpler model. The early kilonova, which is dominated by the photospheric emission, is therefore brighter and peaks at higher frequencies. Moreover, the more efficient thermalization keeps the ejecta hotter for longer, shifting the peak at later times, around $t\sim 2$ hours. The thermalization efficiency drops exponentially for $t\gtrsim 0.1$ day. At late times ($t\gtrsim 1$ day), the advanced model predicts a smaller contribution to the light curves from the energy thermalized in the optically thin layers of the outflow. The material is overall colder, and so is the photosphere (despite its faster recession due to the multi-group opacities). Both the photospheric emission and the contribution to the luminosity from the optically thin layers are therefore suppressed by the composition-dependent thermalization, explaining the steeper late-time decline. The LANL opacity model further reinforces this suppression because of the faster recession of the photosphere, which leads to an earlier drop in the infrared bands.

### v2

@@ 3.3 Effects of thermalization treatment
The thermalization prescription directly enters the nuclear term in Eq. (1). However, nuclear heating remains a subdominant contribution to the dynamics for most of the ejecta evolution. We show this behavior in Fig 8, which displays the instantaneous average contributions to the internal energy of the fluid for the BLh profile under different thermalization and opacity prescriptions. On the timescales relevant for the nucleosynthesis, the effects of the different thermalization schemes on the temperature and density evolution are almost negligible. As a result, for the BLh profile, the final nuclear yields agree within roughly $\sim 10\%$, and the final mass fraction of lanthanides is consistent to within a few percent (see Table 4). The isotopic evolution is similar for both thermalization schemes. Neutron freeze-out occurs slightly earlier (by $\sim 200$ ms) in the simple thermalization case, with no qualitative change in the rest of the isotopic evolution.
@@
At early times, the thermalization efficiency is close to unity for all decay products. Any constant, fractional thermalization factor therefore underestimates the nuclear contribution to the internal energy of the ejecta. The resulting lower temperature of the photosphere explains the dimmer and redder light curves produced by the simple thermalization scheme at $t\!\sim$ hours and the lower bolometric luminosity reported in Table 4. The two models cross between roughly $\sim\!6$ hours and $\sim\!1$ day, when the effective averaged thermalization factor from the online calculations drops to $\langle f_{\text{th}}\rangle\sim 0.5$.
@@ 3.4 Effects of opacity models
The (optical) opacity enters the luminosity term in Eq. (1), and can influence the ejecta dynamics. Consequently, a different opacity model could also indirectly affect the composition evolution. However, since radiation contributes only a subdominant fraction of the energy balance (see Fig. 8), a detailed multigroup treatment of the opacity is not necessary to accurately capture the hydrodynamics. The radiation term becomes relevant only at $t\gtrsim 1$ days, when the expansion is already fully homologous and changes in the composition of the outflow are driven only by nuclear decays. An order-of-magnitude estimate of the gray opacity is therefore sufficient to correctly reproduce the ejecta dynamics and perform reliable nucleosynthesis calculations. The relative difference between the abundances predicted by different opacity models always remain below a few $0.1\%$, and the final mass fraction of lanthanides is essentially unchanged (see Table 4).
@@
Figure 11: Evolution of the gray opacity $\kappa_{\textrm{gray}}$ for a sampling of fluid elements across all angles and depths, from simulations using different opacity models. Top row: Runs using the analytical expression from Wu et al. (2022) (left) or the prescription from Just et al. (2022) (right). Central row: Simulations with Rosseland (left) or Planck (right) gray opacities computed from the LANL data for neodymium as representative of the lanthanides. Bottom row: Rosseland mean opacity from the multigroup models using only neodymium to represent lanthanides (left) or also uranium as representative of actinides (right). The color bar is cut at $\kappa_{\textrm{gray}}=10^{-2}$ cm2g-1 for easier visualization. For each simulation, the top and bottom subpanels show the trajectories of the fluid elements in the density-temperature and density-cumulative lanthanides mass fraction ($X_{\text{lan}}$) planes, respectively. Markers indicate different orders of magnitudes of $X_{\text{lan}}$, with circles for $X_{\text{lan}}<5\times 10^{-5}$, downward triangles for $10^{-5}<X_{\text{lan}}<10^{-3}$, upward triangles for $10^{-3}<X_{\text{lan}}<5\times 10^{-2}$, and stars for $X_{\text{lan}}>5\times 10^{-2}$. For the simulations using only neodymium data, the background shows the single-element mean opacity $\kappa_{\textrm{el}}$ calculated from the LANL tables (Rosseland for GK and Nd300, Planck for GKPl). Similarly, the uranium Rosseland mean opacity is shown in the NdU300 panel.
The same reasoning applies when comparing the multigroup calculations with the J[22] model, employed for the JK run. Although this prescription adapts better to the multigroup results based on neodymium (and uranium) data (simulations Nd300 and NdU300), it still overestimates the opacity in the (equatorial) lanthanide-active ($X_{\text{lan}}\gtrsim 10^{-5}$) regions at temperatures around $T\sim 10^{3}$ K.
@@
The nonanalytical models exhibit non-monotonic opacity evolution in the lanthanide-rich elements. This behavior arises either from the opacity peak around $T\sim 5\times 10^{3}$ K in the LANL-R and LANL cases, or from the activation of the atomic data at $T\simeq 6\times 10^{4}$ K in the LANL-P model (see Fig. 11). As a result, the photosphere is observed to bounce toward the outer regions of the ejecta at $t\sim 1$ hour in the GKPl simulation, and later at $t\sim 5$ hours in the GK, Nd300, and NdU300 runs. The photosphere location predicted by the multigroup calculations always lies between those obtained with LANL-R and LANL-P, which are known to respectively under- and overestimate the effective gray opacity.
@@
In Fig. 13, we examine the impact of the opacity model on the kilonova light curves. The colder photosphere predicted by W[22] with respect to the LANL model results in a suppressed emission at $t\lesssim 1$ hour (in agreement with the bolometric luminosities reported in Table 4), with a stronger suppression at higher frequencies that makes the early-time signal appear redder. Nevertheless, the photospheric luminosity and the optical-band magnitudes remain broadly comparable across models, as the larger photospheric radius partially compensates for the lower emissivity of the colder photosphere. At later times, the artificially high constant gray opacity causes the ejecta to become completely optically thin only at comparatively late stages. Therefore, the photospheric contribution to the luminosity persists for longer, thus delaying the decay of the infrared emission at $t\gtrsim 1$ day.
@@
Figure 13: Predicted AB apparent magnitudes in the Gemini $u$, $g$, $i$ and $K_{s}$ bands for an observer at a polar angle of $30$ degrees and a distance of 40 Mpc. Comparison between a series of opacity models applied to the BLh profile. Left: Frequency- and composition-dependent opacities based on atomic calculations for neodymium and uranium (solid lines), versus the Just et al. (2022) prescription (dashed) and the simple Wu et al. (2022) opacity (dotted). Center: Opacities based on the neodymium atomic calculations, multigroup (solid lines) versus gray models (Rosseland or Planck, dashed and dotted). Right: Frequency-dependent LANL opacities including uranium (solid lines) or only neodymium (dashed) compared against the analytical Just et al. (2022) model (dotted).
@@
As discussed above, J[22] performs better than W[22], but still overestimates the opacity of the lanthanide-rich ejecta. As a result, the light curves from J[22] more closely resemble those produced with the LANL models than those from W[22], but they still exhibit a lower peak bolometric luminosity and a delayed shut-off of the photospheric emission and thus a more prolonged signal.
@@
The central panel of Fig. 13 compares simulations using gray (GK and GKPl) and multigroup (Nd300) opacities based on neodymium atomic calculations. As expected, the Rosseland mean opacity performs better in the light-curve calculations at early times, when the ejecta is still optically thick. At later times, the Planck opacity provides a better approximation for the kilonova emission, especially in the higher-frequency bands. The two gray approximations bracket the actual effective opacity, and therefore generally bracket the multigroup light curves across all filters over most of the evolution. In particular, the Planck mean misses the early blue/UV peak at $t\sim 0.1$ days, whereas the Rosseland opacity predicts too bright and delayed magnitude peaks, especially in the higher-frequency filters. The peak bolometric luminosity is consequently enhanced and delayed by the LANL-R model (see Table 4). At late times ($t\gtrsim 10$ days), the Planck opacity performs worse again because of the delayed recession of the photosphere in a more opaque material (see Fig. 12). This effect dominates over the direct suppression of the emergent flux (e.g. Eqs. (3) and (19)), leading to an inversion of the hierarchy of the models in the red filter at $t\gtrsim 5$ days.
@@
The right panel of Fig. 13 shows that J[22] reproduces multigroup results based on neodymium data more accurately than our gray, atomic-physics-based models. This result highlights the necessity of employing frequency-dependent transport when using detailed atomic opacities, as compressing the information from line-resolved calculations into effective gray averages can be misleading to the point of being outperformed by analytical fitting models.
@@
The original work of Kasen et al. (2017) employed atomic calculations only up to the lanthanides, which at least partially explains the disagreement with our most sophisticated opacity model that also includes uranium (NdU300). Fontes et al. (2020) demonstrated differences between semi-relativistic and fully relativistic neodymium opacities. Additional improvements in opacity modeling may be achieved by replacing theoretically calculated energy levels with NIST-calibrated values. Incorporating these calibrated energies may affects the corresponding light curves and spectra at late times.
@@
When only neodymium is incorporated, the effective opacities predicted by the frequency-dependent simulations lie between the Rosseland and Planck mean values, but are qualitatively closer to the latter (see Fig. 11). Correspondingly, the J[22] approximation produces light curves more similar to LANL-P, particularly at late times, when the ejecta becomes optically thin. However, the analytical prescription fails to capture the drop in effective opacity for the $r$-process fluid elements at $T\sim 2-3\times 10^{4}$ K. As a result, it misses the luminosity peak at $t\sim 1-3$ hours, caused by regions of the outflow that are hotter than their surroundings due to early ($t\sim 1$ s) decays of freshly synthesized $r$-process elements (Bernuzzi et al., 2025). This missing peak, already visible in the original publication by Just et al. (2022), is more relevant for trajectories with higher lanthanide fractions.
@@ 3.5 Impact of a polar jet
In Fig. 14, we show the evolution of the radial velocity and temperature, each rescaled to the interval $[0,1]$ over the plotted time range, for a few representative fluid elements on the timescales relevant for the GRB. The activation of the thermal bomb generates a shock visible in the radial velocity of the innermost mass shells. In the most polar regions, the same fluid elements also experience a rapid increase in temperature. The impact of the jet becomes progressively weaker at larger polar angles, where the deposited energy is smaller. For outer mass shells, the response is delayed, reflecting the finite propagation time of the shock through the ejecta.
Figure 14: Radial velocity (top) and temperature (bottom) for a set of fluid elements from the GK simulation including the polar jet. Both quantities are rescaled to the interval $[0,1]$ for visualization purposes. Colors and styles correspond to different angular sections or mass-shell indices, respectively. The vertical black lines indicate the duration of the jet.
As already commented in Magistrelli et al. (2024), the impact of the jet on the globally averaged nucleosynthesis is negligible. Figure 15 shows the nuclear yields from the polar regions ($\theta\lesssim 15$ degrees) of the BLh ejecta in the structured-jet and no-jet configurations. Because this angular cut excludes the neutron-rich equatorial outflow, the resulting nucleosynthesis is dominated by the production of elements up to the first peak. The nuclear yields are essentially unchanged by the introduction of the jet, with typical discrepancies on the order of a few tens of percent or less.
@@ 3.6 Best model comparison
In this section, we summarize the combined effects of all the improvements that we implemented on top of the code from Wu et al. (2022). We compare our most advanced setup (NdU300), featuring the in-situ NN, detailed thermalization, multigroup radiative transfer, and atomic-physics based opacities, against the results obtained using the 2D ray-by-ray extension of original version of the code (Apr2_BLh). For the latter, we also implement the generalized fits for the nuclear power discussed in Sec. 2.3.
As already mentioned in the previous sections, the improved thermalization and opacity prescriptions do not significantly affect the composition evolution of the ejecta. Slightly larger discrepancies appear in the nucleosynthesis yields and abundances evolution with respect to those described in Sec. 3.1, but the physical considerations remain the same. The introduction of the in-situ NN is the dominant source of variation in the nucleosynthesis results, and a detailed description of the hydrodynamic evolution of the ejecta is essential for reliable nucleosynthesis predictions. In particular, assuming homologous expansion is not safe on $r$-process timescales, leading to deviations greater than unity for most mass numbers and a shifted third $r$-process peaks (see the green and red curves in the bottom panels of Fig. 3). On the other hand, calculating the nuclear power on the fly and including its back-reaction on the hydrodynamics has a comparatively minor impact, typically at the $\lesssim 20\%$ level (see the orange lines in the same panels). For the purpose of studying element formation, the hydrodynamics of the ejecta can therefore be reproduced well enough by performing simulations with accurately calibrated nuclear power fits. A post-processing NN can then recover the results of the online calculations within a few $10\%$, provided that the set of tracers is sufficiently large to properly resolve both the dynamics and the initial composition distribution of the ejecta.
@@
Figure 17: Late-time ($t\simeq 3$ days) ratios between the mass-weighted histograms of the density, temperature and instantaneous thermalized heating rate obtained from the BLh runs using the in-situ NN and our most advanced thermalization and opacity prescriptions (NdU300), or from the 2D ray-by-ray extension of the configuration of Wu et al. (2022) (Apr2_BLh). The left [right] column shows the ratios as functions of the polar angle [initial electron fraction]. Bins containing less than $10^{-5}M_{\odot}$ in both simulations are plotted with increasing transparency for decreasing mass content (considering the run with most mass in that bin). Thus, the color encodes the value of the ratio in $\Delta M$, while the saturation reflects the highest between the two original values $\Delta M$ (cf. Fig. 5)
@@
The differences in temperature evolution and thermalized heating rates have an immediate impact on the predicted light curves. Figure 18 displays the kilonovae associated with the Apr2_BLh and NdU300 simulations, both showing the characteristic blue/UV peak at $t\sim 1$ hour and the later red emission. At early times ($t\lesssim 0.1$ day), the average thermalization efficiency is comparable between the two simulations, and higher in the DT08 scheme. At the same time, the multigroup opacities drive a faster recession of the photosphere. Both effects contribute to a hotter photosphere than in the simpler model. The early kilonova, which is dominated by the photospheric emission, is therefore brighter and peaks at higher frequencies. Moreover, the more efficient thermalization keeps the ejecta hotter for longer, shifting the peak at later times, around $t\sim 2$ hours. The thermalization efficiency drops exponentially for $t\gtrsim 0.1$ day. At late times ($t\gtrsim 1$ day), the advanced model predicts a smaller contribution to the light curves from the energy thermalized in the optically thin layers of the outflow. The material is overall colder, and so is the photosphere (despite its faster recession due to the multigroup opacities). Both the photospheric emission and the contribution to the luminosity from the optically thin layers are therefore suppressed by the composition-dependent thermalization, explaining the steeper late-time decline. The LANL opacity model further reinforces this suppression because of the faster recession of the photosphere, which leads to an earlier drop in the infrared bands.